TFY4340 Mesoscopic Physics Spring 2010

Course information

Updated: January 14, 2010.

Course webpage:

http://web.phys.ntnu.no/~stovneng/TFY4340_2010/tfy4340.htm

Lecturer: Jon Andreas Støvneng Department of physics, NTNU, 7491 Trondheim Office: E5-130 Phone: 73 59 36 63 Email: jon.stovneng@ntnu.no

Timetable:

Day  Hours  Room Activity
Monday 09.15 – 10.00 E5-111 Lecture
Monday 10.15 – 12.00 E5-111 Exercise
Tuesday 08.15 – 10.00 E5-111 Lectures

Comment: Room E5-111 is in the Theory section of Department of physics, on level 5, E corridor.

Syllabus:

• To a first approximation, lectured material and given exercises.

• A more precise definition will come later.

Books etc: Parts of the books by Heinzel and Datta will cover most of the syllabus.

Evaluation: According to the official NTNU course information, there is supposed to be an oral examination on May 28 2010, counting 100 %. In practice, I would like to discuss this with you students.

Department of physics, January 14, 2010. Jon Andreas Støvneng